# Structural characterizations and Obstructions to Planar-Rips complexes

### v1

@@ 5. Obstructions to planar-Rips complexes and correlations with Unit Disk Graphs
Given any finite set $X\subset\mathbb{R}^{2}$, the planar-Rips complex of $X$ is same as the clique complex of the unit disk graph, having the vertex set as $X$.
For simplicity, we shall fix $r=1$ given the scale-invariant nature of our study. Hence, for $X\subset\mathbb{R}^{2}$, we shall denote the unit disk graph on the vertex set $X$ by $\mathcal{R}^{1}(X)$, which denotes the $1$-skeleton of planar-Rips complex.
Consequently, there is a one-to-one correspondence, implying that both the categories are isomorphic. This correspondence facilitates the integration of both studies, enabling the translation of the statements and results in the category of planar-Rips complexes into corresponding analogs for the category of unit disk graphs, or vice-versa.
In the preceding sections, a complete characterization of planar-Rips complexes with $n$-dimensional weak-pseudomanifold structure and minimal $2$-cycle structure has been outlined, including a partial characterization of two-dimensional, pure, and closed planar-Rips complexes. These structures correspond to unit disk graphs having all maximal cliques of size $(n+1)$, and every clique of size $n$ being a part of exactly two maximal cliques. Furthermore, a relation can be drawn for unit disk graphs corresponding to the property of ’strongly connected’ness of planar-Rips complexes. To make these characterizations readily accessible in terms of unit disk graphs, we redefine some graph analogs corresponding to pseudomanifolds, weak-pseudomanfolds and closed simplicial complexes:
@@
    A clique is said to be maximal if it is not contained in a clique of larger size.
@@
  3. (c)
    A pure graph (with clique number $n$) is said to be strongly connected if for any two maximal cliques $\sigma\text{ and }\gamma$, there exists a sequence of maximal cliques, say $\lambda_{1},\lambda_{2},\dots\lambda_{k},$ such that $\lambda_{1}=\sigma,\ \lambda_{k}=\gamma$ and $\lambda_{i}\cap\lambda_{i+1}$ is a clique of size $(n-1)$ for all $11\leq i\leq k-$ 1.
  4. (d)
    A pure graph (with clique number $n$) is said to be closed if every clique of size ($n-1$) is contained in at least two maximal cliques.
@@
5.1. Graph of cross-polytopes:
Reframe these statements for my research paper : ”The class of planar-Rips complexes have an ability to be characterized by the forbidden planar-Rips structures. This section is specifically streamlined towards exploring this fascinating feature of class of planar-Rips complexes.”
The graph of cross-polytopes appears most often as the induced subcomplex in pure and strongly connected unit disk graphs (See Corollary 5.4). The graph of $1$-skeleton of the boundary complex of an $n$-dimensional cross-polytope is isomorphic to a complete $n$-partite graph with each independent set of size $2$. For instance, the graph of octahedron is isomorphic to the complete tripartite graph $K_{2,2,2}$ (See Figure 13.) Henceforth, we shall also refer to these graph as complete bi-independent $n$-partite graph. We will denote this graph by $K_{n\times 2}$ as defined by Brouwer et al. [10, p. 478], to mean the complete $n$-partite graph $K_{\underbrace{2,2,\dots,2}_{n\text{ times}}}$.
Figure 13. Complete bi-independent $3$-partite graph $K_{2,2,2}$
Now we will state some of the corollaries which are natural graph analogs to the theorems proved in the previous sections. These corollaries reveal interesting properties of unit disk graph structures, demonstrating the power and generality of our approach, and open up new avenues for our further study on forbidden structures for planar-Rips complexes and unit disk graphs:
@@ 5.1. Graph of cross-polytopes:
Let $G$ be a pure and closed unit disk graph with $\omega(G)=3$ then
@@
These simply follow from Theorem 4.7, Proposition 4.4, Remark 4.13, and Proposition 4.14 by considering the graph of $1$-skeleton of two-dimensional pure and closed planar-Rips complexes. ∎
Remark 5.5.
For the above theorem, it is easy to verify that a graph $G$ is pure and closed with clique number $\omega(G)=3$ if and only if $G$ is connected with $\omega(G)=3$, as stated in D, where every edge has a degree of at least $2$.
@@
This simply follows from Theorem 3.18 by considering the graph of $1$-skeleton of planar-Rips complexes with weak-pseudomanifold structure. ∎
5.2. Obstructions to Planar-Rips complexes
Given a hereditary property $\mathcal{P}$ in the category of simplicial complexes or the category of graphs, any complex or a graph that does not satisfy $\mathcal{P}$ further contains a minimal induced substructure, say $F$, that does not satisfy $\mathcal{P}$. The minimality is defined such that all the induced substructures of $F$ do satisfy $\mathcal{P}$. The existence of such an $F$ can be easily verified with repetitive vertex deletions. In case of simplicial complexes, these structures are referred to as obstructions to the corresponding property. We will now define obstructions to planar-Rips complexes, where the property $\mathcal{P}$ denotes whether the simplicial complex admits a planar-Rips structure.
@@ 5.2. Obstructions to Planar-Rips complexes
An obstruction to planar-Rips complex is a simplicial complex that does not admit planar-Rips structure, all of whose proper induced subcomplexes do.
We say a graph $G$ admits a unit disk graph structure if there exists an injective map $f:V(G)\to\mathbb{R}^{2}$ such that $G$ is isomorphic to the unit disk graph $\mathcal{R}^{1}(f(V))$. In other words, $uv$ is an edge if and only if the distance between $f(u)$ and $f(v)$ is less than $1$. In case of unit disk graphs, the analog of obstructions are minimal forbidden induced subgraphs for unit disk graphs, referred to as minimal non unit disk graphs [4].
@@
There is a close relationship between the minimal non-UDGs and the obstructions to planar-Rips complexes. The following is a straightforward inference:
Claim.
The clique complex of a minimal forbidden induced subgraph for unit disk graph is an obstruction to planar-Rips complex.
However, it is not true that the $1$-skeleton of every obstruction to planar-Rips complex is a minimal forbidden induced subgraph for unit disk graph. For instance, the boundary complex of all $n$-dimensional simplices are obstructions to planar-Rips complex. Considering that these are also the obstructions to flag complexes, therefore, we shall restrict only to the class of flag complexes to have the one-to-one correspondence as follows:
@@
Let $K$ be a flag complex, then $K$ is an obstruction to planar-Rips complex if and only if the graph of $1$-skeleton $K^{(1)}$ is a minimal non-UDG.
@@
Therefore, it follows that the clique complex of a minimal non-UDG is an obstruction to planar-Rips complex. The literature on minimal non-UDGs is surprisingly sparse despite active research in unit disk graphs. The following theorem provides a concise summary of currently known minimal non-UDGs from [4, 5], providing us with a few obstructions to planar-Rips complex.
[…]
Figure 14. Known minimal non unit disk graphs with $5$, $6$ and $7$ vertices (Source: [5, p. 2])
@@
The following graphs are minimal non-UDGs. More precisely, the clique complex of these graphs are obstructions to planar-Rips complex.
@@
Note that in the above theorem, the graph $C^{*}_{2k}$ denotes the graph obtained by adding two $k$-cliques $\overline{U}$ and $\overline{V}$ to the cyclic graph $C_{2k}$, where $U$ and $V$ are the two maximal independent sets in $C_{2k}$, each of size $k$. Furthermore, it follows that there is an obstruction to planar-Rips complex in every odd dimension $d$.
In the realm of topology, the classification problem, namely, classification of structures like manifolds, pseudomanifolds and weak-pseudomanifolds up to homeomorphism (or homotopy) has been extensively studied and still remains a hot topic. Indeed, these structures have been largely classified in the literature. Following Theorem 2.3, Theorem 4.7 and Lemma 3.6, we have a large class of forbidden induced structures for planar-Rips complexes and unit disk graphs. For instance, the result by Adamszek in Theorem 3.3 implies that no manifold other than the cross-polytopal spheres $\mathcal{O}_{n}$ admits a planar-Rips structure. This makes a pavement for extracting a wide spectrum of obstructions from these forbidden structures. Consider the following examples.
@@
Let $K$ be the complex $C_{5}*\mathbb{S}^{0}$, a flag triangulation of sphere $\mathbb{S}^{2}$ where $C_{5}$ is a cycle of length $5$. The complex $K$ being a $2$-dimensional pseudomanifold does not admit a planar-Rips structure as follows from Theorem 3.7. But is the complex $K$ an obstruction to planar-Rips complex? The $1$-skeleton of $K$ can be identified as the complement of the graph $\overline{K_{2}+C_{5}}$ which is a minimal non-UDG as shown in [5, Theorem 4.1]. Therefore, the complex $K$ is indeed an obstruction to the planar-Rips complex.
@@
Figure 15. Representation of minimal flag triangulation of $\mathbb{R}P^{2}$ in terms of its $1$-skeleton (on the left) that does not admit a planar-Rips structure. Its induced subcomplex ${\left\langle{ABCDEFG}\right\rangle}$ (on the right) is an obstruction to planar-Rips complex
@@
Consider the minimal triangulation of the projective space $\mathbb{R}P^{2}$ with $11$ vertices as shown in Figure 15(a) [12, Example 3.1]. We know that this complex does not admit a planar-Rips structure since its fundamental group is not free. Hence, there exists at least one obstruction to planar-Rips complex that is embedded inside this triangulation. The complex being non-orientable and with a torsion in its homology group, it appears to be a ‘highly forbidden’ structure for a planar-Rips complex, meaning that an embedded obstruction should be repetitive in its structure. With this said, we consider the induced subcomplex $K={\left\langle{ABCDEFG}\right\rangle}$ and consequently prove in the next theorem that the graph of $1$-skeleton of $K$ is a minimal non-UDG. In other words, $K$ is indeed an obstruction to planar-Rips complex.
@@
The graph in Figure 15(b) is a minimal non-UDG, or equivalently, its clique complex is an obstruction to planar-Rips complex.
@@
Let $K={\left\langle{ABCDEFG}\right\rangle}$ denote the clique complex of the graph in Figure 15(b). We first show that $K$ does not admit a planar-Rips structure. On contrary, suppose it admits a planar-Rips structure; Without loss of generality, let $\mathcal{R}(X)$ denote the planar-Rips structure, where $X=\left\{A,B,C,D,E,F,G\right\}\subset\mathbb{R}^{2}$. Note that the fundamental group $\pi_{1}(\mathcal{R}(X))$ is non-trivial.
@@
Consider the line $L_{CB}$ and the perpendicular bisectors $L_{C|B},L_{D|C}$, and $L_{C|E}$. Firstly, $E$ cannot be on the line $L_{CB}$ since $E$ is not adjacent to $B$, and $C$ is not adjacent to $F$. Likewise, $D$ cannot be on the line $L_{CB}$. Without loss of generality, assume $E$ is on the right side of the line $L_{CB}$ i.e. $E\in L^{-}_{CB}$. Considering the adjacency between the vertices, we have the following observations:
  1. (a).
@@
With this arrangement of vertices, the segment $\overline{CB}$ lies on the left side of every line joining any two adjacent vertices of polygon $DCEFBG$ (in counterclockwise order). See Figure 16 for an illustration. It follows that $\overline{CB}\in\mathrm{conv}\left\{C,D,E\right\}\cup\mathrm{conv}\left\{D,E,F,B,G\right\}$. ∎
@@
Additionally, it can be easily verified that every induced subcomplex of $K$ admits a planar-Rips structure. Therefore, $K$ is an obstruction to planar-Rips complex. ∎
@@ 6. Concluding remarks
In a combinatorial framework, the topological methods stand out for their robust ability to uncover locally governed features that are pivotal for reconstructing the global structure. The methods and results presented in this paper highlight the effectiveness of topological and geometric approaches towards characterizing planar-Rips complexes by capturing the local constraints. In particular, the concept of obstructions to planar-Rips complexes has been introduced, a tool that can aid in recognizing forbidden planar-Rips structures. Notably, this concept has a natural generalization to Vietoris-Rips complexes inside any metric space. It is also worth noting that the challenge of determining the obstructions to planar-Rips complexes can be reformulated into the following problem:
@@
Given a flag complex $K$ that does not admit a planar-Rips structure, devise methodologies to determine the induced subcomplex(es) of $K$ that are obstructions to planar-Rips complex.
Developing algorithms for recognizing forbidden planar-Rips complexes presents another significant avenue for research. It would be interesting to know how far this approach can be developed. Moreover, our work progresses towards proving the conjecture [2] that every planar-Rips complex is homotopy equivalent to a wedge of spheres.

### v2

@@ 5. Correlations with unit disk graphs
Given any finite set $X\subset\mathbb{R}^{2}$, the planar-Rips complex of $X$ is same as the clique complex of the disk graph, having the vertex set as $X$.
For simplicity, we shall fix $\delta=1$ given the scale-invariant nature of our study. Henceforth, we shall denote the unit disk graph on the vertex set $X\subset\mathbb{R}^{2}$ by $\mathcal{R}^{1}(X)$, the $1$-skeleton of planar-Rips complex.
Consequently, there is a one-to-one correspondence, implying that both the categories are isomorphic. This correspondence facilitates the integration of both studies, enabling the translation of the statements and results in the category of planar-Rips complexes into corresponding analogs for the category of unit disk graphs, or vice-versa. We say a graph $G$ admits a unit disk graph structure if there exists an injective map $f\colon V(G)\to\mathbb{R}^{2}$ such that $G$ is isomorphic to the unit disk graph $\mathcal{R}^{1}(f(V))$ (we will denote this by $\mathcal{D}G$). In other words, $uv$ is an edge if and only if the distance between $f(u)$ and $f(v)$ is less than $1$ in $\mathcal{D}G$.
@@
    A clique is said to be maximal if it is not contained in a clique of larger cardinality. We will denote the cardinality of the largest clique in a graph $G$ by $\omega(G)$.
@@
5.1. Graph of cross-polytopes.
The graph of cross-polytopes appears most often as the induced subgraph in pure unit disk graphs (See Corollary 5.4). The graph of $1$-skeleton of the boundary complex of an $n$-dimensional cross-polytope is isomorphic to a complete $n$-partite graph with each independent set of size $2$. For instance, the graph of octahedron is isomorphic to the complete tripartite graph $K_{2,2,2}$ (See Figure 14.) Henceforth, we shall also refer to these graph as complete bi-independent $n$-partite graph. We will denote this graph by $K_{n\times 2}$ as defined by Brouwer et al. [11, p. 478], to mean the complete $n$-partite graph $K_{\underbrace{2,2,\dots,2}_{n\text{ times}}}$.
Figure 14. Complete bi-independent $n$-partite graph $K_{n\times 2}$
In the preceding sections, a complete characterization of planar-Rips complexes with $n$-dimensional weak-pseudomanifold structure and minimal $2$-cycle structure have been outlined, including a partial characterization of two-dimensional, pure, and closed planar-Rips complexes. These structures correspond to unit disk graphs having all maximal cliques of size $(n+1)$, and every clique of size $n$ contained in exactly two maximal cliques. Now we will state some of the corollaries which are natural graph analogs to these theorems proved in the previous sections.
@@
Let $G$ be a connected graph having clique number $\omega(G)=3$ and every edge is contained in at least two maximal cliques. If $G$ admits a unit disk graph structure, then
@@
Suppose $\mathcal{R}^{1}(X)$ is the unit disk graph representation of $G$ for some $X\subset\mathbb{R}^{2}$. Since $\omega(G)=3$, the planar-Rips complex $\mathcal{R}(X)$ is two-dimensional. Further, $\mathcal{R}(X)$ is pure and closed as follows from the hypothesis that $G$ is connected, and every edge is contained in at least two maximal cliques. Now the conclusion simply follows from Theorem 4.7, Proposition 4.4, Remark 4.13, and Proposition 4.14, where the graph of $1$-skeleton of octahedron is $K_{2,2,2}$. ∎
@@
This simply follows from Theorem 3.18 by considering the graph of $1$-skeleton of planar-Rips complexes, having weak-pseudomanifold structure. ∎
5.2. Obstructions of Planar-Rips complexes
Given a hereditary property $\mathcal{P}$ in the category of simplicial complexes or the category of graphs, any complex or a graph that does not satisfy $\mathcal{P}$ further contains a minimal induced substructure, say $F$, that does not satisfy $\mathcal{P}$. The minimality is defined such that all the induced substructures of $F$ do satisfy $\mathcal{P}$. The existence of such an $F$ can be easily verified with repetitive vertex deletions. In case of simplicial complexes, these structures are referred to as obstructions to the corresponding property. We will now define obstructions of planar-Rips complexes, where the property $\mathcal{P}$ denotes whether a simplicial complex admits a planar-Rips structure.
@@ 5.2. Obstructions of Planar-Rips complexes
An obstruction of planar-Rips complex is a simplicial complex that does not admit planar-Rips structure, all of whose proper induced subcomplexes do.
In case of unit disk graphs, the analog of obstructions is minimal forbidden induced subgraphs for unit disk graphs, referred to as minimal non unit disk graphs [20].
@@
There is a close relationship between the minimal non-UDGs and the obstructions of planar-Rips complexes. The following observation is a straightforward inference:
Observation.
The clique complex of a minimal forbidden induced subgraph for unit disk graph is an obstruction of planar-Rips complex.
However, it is not true that the $1$-skeleton of every obstruction of planar-Rips complex is a minimal forbidden induced subgraph for unit disk graph. For instance, the boundary complex of all $n$-dimensional simplices are obstructions of planar-Rips complex. Considering that these are also the obstructions of flag complexes, therefore, we shall restrict only to the class of flag complexes to have the one-to-one correspondence as follows:
@@
Let $K$ be a flag complex, then $K$ is an obstruction of planar-Rips complex if and only if the graph of $1$-skeleton $K^{(1)}$ is a minimal non-UDG.
@@
Therefore, it follows that the clique complex of a minimal non-UDG is an obstruction of planar-Rips complex. The literature on minimal non-UDGs is surprisingly sparse despite active research in unit disk graphs. The following theorem from [20, 5] provides a concise summary of currently known obstructions of planar-Rips complexes in terms of minimal non-UDGs.
$G_{2}$$G_{7}$$G_{6}$$G_{5}$$G_{4}$$G_{3}$$G_{1}$
Figure 15. Known minimal non unit disk graphs with $\leq 7$ vertices (Source: [5, p. 2])
@@
The following graphs are minimal non-UDGs. More precisely, the clique complex of these graphs are obstructions of planar-Rips complex.
@@
Note that in the above theorem, the graph $C^{*}_{2k}$ denotes the graph obtained by adding two $k$-cliques $\overline{U}$ and $\overline{V}$ to the cyclic graph $C_{2k}$, where $U$ and $V$ are the two maximal independent sets in $C_{2k}$, each of size $k$. Furthermore, it follows that there is an obstruction of planar-Rips complex in every odd dimension $d$.
In the realm of topology, the classification problem, namely, classification of structures like manifolds, pseudomanifolds and weak-pseudomanifolds up to homeomorphism (or homotopy) has been extensively studied and still remains a hot topic. Indeed, these structures have been largely classified in the literature [7, 6, 12, 26]. Following Theorem 2.3, Theorem 4.7 and Lemma 3.6, we have a large class of forbidden induced structures for planar-Rips complexes and unit disk graphs. For instance, the result by Adamaszek et al. in Theorem 3.3 implies that no manifold other than the cross-polytopal spheres $\mathcal{O}_{n}$ admits a planar-Rips structure. This makes a pavement for extracting a wide spectrum of obstructions from these forbidden structures. Consider the following examples.
@@
Let $K$ be the complex $C_{5}*\mathbb{S}^{0}$, a flag triangulation of sphere $\mathbb{S}^{2}$ where $C_{5}$ is a cycle of length $5$. The complex $K$ being a $2$-dimensional pseudomanifold does not admit a planar-Rips structure as follows from Theorem 3.7. To determine whether $K$ is an obstruction of a planar-Rips complex, we note that the $1$-skeleton of $K$ can be identified as the complement of the graph $\overline{K_{2}+C_{5}}$ which is a minimal non-UDG as shown in [5, Theorem 4.1]. Therefore, the complex $K$ is indeed an obstruction of planar-Rips complex.
@@
Figure 16. Representation of minimal flag triangulation of $\mathbb{R}P^{2}$ in terms of its $1$-skeleton (on the left) that does not admit a planar-Rips structure. Its induced subcomplex $\left\langle{ABCDEFG}\right\rangle$} (on the right) is an obstruction of planar-Rips complex
@@
Consider the minimal flag triangulation of the projective space $\mathbb{R}P^{2}$ with $11$ vertices as shown in Figure 16(a) [8, Example 3.1]. We know that this complex cannot be realized as a planar-Rips complex since its fundamental group is not free. Hence, there exists at least one obstruction of planar-Rips complex that is embedded inside this triangulation. The complex being non-orientable and with a torsion in its homology group, it appears to be a ‘highly forbidden’ structure for a planar-Rips complex, suggesting that any embedded obstruction within it is likely to exhibit repetitive structural patterns. With this in mind, we consider the induced subcomplex $K={\left\langle{ABCDEFG}\right\rangle}$, and consequently use topological methods to prove in the next theorem that $K$ is indeed an obstruction of planar-Rips complex. In other words, the graph of $1$-skeleton of $K$ is a minimal non-UDG.
@@
The graph in Figure 16(b) is a minimal non-UDG, or equivalently, its clique complex is an obstruction of planar-Rips complex.
@@
Let $K={\left\langle{ABCDEFG}\right\rangle}$ denote the clique complex of the graph in Figure 16(b). We first show that $K$ does not admit a planar-Rips structure. On contrary, suppose it admits a planar-Rips structure; Without loss of generality, let $\mathcal{R}(X)$ denote the planar-Rips structure, where $X=\left\{A,B,C,D,E,F,G\right\}\subset\mathbb{R}^{2}$. Note that the fundamental group $\pi_{1}(\mathcal{R}(X))$ is non-trivial since the cycle $\left\langle{BCEF}\right\rangle$} at the vertex $B$ is not contractible in $\mathcal{R}(X)$.
@@
Consider the line $L_{CB}$ and the perpendicular bisectors $L_{C|B},L_{D|C}$, and $L_{C|E}$. Firstly, $E$ cannot be on the line $L_{CB}$ since $E$ is not adjacent to $B$, and $C$ is not adjacent to $F$. Likewise, $D$ cannot be on the line $L_{CB}$. Without loss of generality, assume $E$ is on the right side of the line $L_{CB}$ i.e., $E\in L^{-}_{CB}$. Considering the adjacency between the vertices, we have the following observations:
  1. (a)
@@
With this arrangement of vertices, the segment $\overline{CB}$ lies on the left side of every line joining any two adjacent vertices of polygon $DCEFBG$ (in counterclockwise order). See Figure 17 for an illustration. It follows that $\overline{CB}\in\mathrm{conv}\left\{C,D,E\right\}\cup\mathrm{conv}\left\{D,E,F,B,G\right\}$. This proves the claim. ∎
@@
Additionally, it can be easily verified that every induced subcomplex of $K$ admits a planar-Rips structure. Therefore, $K$ is an obstruction of planar-Rips complex. ∎
@@ 6. Concluding remarks
The methods and results presented in this paper highlight the effectiveness of topological and geometric approaches towards characterizing planar-Rips complexes. We have successfully obtained a complete classification of planar-Rips complexes that have pseudomanifold and weak-pseudomanifold structure, demonstrating that they are isomorphic to iterated wedge of crosspolytopal spheres. Furthermore, we characterized the two-dimensional pure and closed simplicial complexes, identifying the octahedron as the only minimal 2-cycle. Perhaps, these structures have homotopy type of wedge sum of spheres.
From an application standpoint, we have explored the isomorphism between the category of planar-Rips complexes and unit disk graphs, providing a characterization of a large class of disk graphs with uniform clique structure. In particular, the concept of obstructions in planar-Rips complexes has been introduced, a tool that offers a new framework for recognizing and classifying these structures. Notably, this concept has a natural generalization to Vietoris-Rips complexes inside any metric space. Following are some of the problems that remain open:
  1. (i)
    While our findings support Adamaszek’s conjecture [3] that the planar-Rips complexes are homotopy equivalent to wedge of spheres, the problem still remains unsolved.
  2. (ii)
    The minimal 2-cycle isomorphism to the boundary of the 2-crosspolytope invites further exploration in higher dimensions. Perhaps, this can be the final piece to proving the above conjecture.
It is also worth noting that the challenge of determining the obstructions of planar-Rips complexes can be reformulated into the following problem:
@@
Given a flag complex $K$ that does not admit a planar-Rips structure, devise methodologies to determine the induced subcomplex(es) of $K$ that are obstructions of planar-Rips complex.
Developing algorithms for recognizing forbidden planar-Rips complexes presents another significant avenue for research. It would be interesting to know how far this approach can be developed.